# A Family of Convex Models to Achieve Fairness through Dispersion Control

### v2

@@ IV Multi-Agent Assignment Problem
Although the formulation in (17) is presented as an MILP, its constraint matrix is the node–arc incidence matrix of a bipartite graph, which is totally unimodular. Consequently, the linear programming relaxation has integral extreme points, and solving it via any linear programming solver yields the optimal integer solution [24]. Equivalently, one can apply the Hungarian algorithm directly to obtain the optimum in $O(|\mathcal{J}|^{3})$ time [25]. In contrast, a natural generalization in which each task carries a weight and each agent has a capacity constraint becomes NP‑hard (the generalized assignment problem) [26]. In this work, we restrict attention to the uncapacitated variant in (17), focusing on illustrating how $(\varepsilon,p)$‑fairness models control the dispersion of the processing times $t_{a}$ across agents in the simplest assignment problem, rather than a general variant of the MAAP.
IV-A Fairness-Constrained Multi-Agent Assignment Problem
@@
Although the problem in (17) is solvable in polynomial time, adding (18) makes the resulting problem NP‑complete. Indeed, for two agents with $\varepsilon=1$, it reduces to the classical partition problem [27] that is known to be strongly NP-complete. As noted in Section II, when $p<\infty$, (18) is a convex constraint—specifically, an $\ell_{p}$‑norm conic constraint (for $p=2$, an $\ell_{2}$‑norm cone is referred to as a second‑order cone). Hence, the fairness‑constrained MAAP is a mixed‑integer convex program for $p<\infty$. When $p=\infty$, applying the reduction in (7) to (18) yields a mixed‑integer linear program that can be solved directly using off‑the‑shelf commercial or open‑source mixed‑integer linear programming solvers. For mixed‑integer convex optimization problems ($2\leqslant p<\infty$), constraint (18) can be lazily outer approximated by linear inequalities [28], enabling MILP solvers to obtain the solution. In the next section, we briefly present the outer approximation procedure for a general $\ell_{p}$-norm cone. We remark that the outer approximation algorithm is well-understood in the literature, and the presentation in the next section is only intended to make the article self-contained for the readers.
IV-B Outer Approximation of $\ell_{p}$-Norm Cone
We start with the definition of an $\ell_{p}$-norm cone [29] of dimension $(n+1)$ as
@@ IV-B Outer Approximation of $\ell_{p}$-Norm Cone
Note that the fairness constraint in (18) for any $p\geqslant 2$ can be represented in the form (19) by introducing an auxiliary variable. It is known in the literature that when convex cones are outer approximated by linear inequalities, it is both theoretically stronger and computationally more efficient to decompose a high-dimensional cone into several lower-dimensional cones and approximate each of those individually, rather than approximating the high-dimensional cone directly [30]. Converting a high-dimensional conic constraint into many low-dimensional cones is called “cone disaggregation” [30]. To that end, we present the steps to disaggregate the $(n+1)$-dimensional $\ell_{p}$-norm cone in (19) into $n$ three-dimensional cones plus a single linear constraint:
@@
In the last equivalence, each constraint $Y_{i}^{\tfrac{1}{p}}\,z^{1-\tfrac{1}{p}}\geqslant y_{i}$ is convex for $2\leqslant p<\infty$ and in fact defines a three-dimensional power cone [29]:
@@
V Computational Results
This section provides empirical validation of the theoretical results in Section III using the formulations from Section IV. All experiments were run on 50 randomly generated instances of the MAAP, each with 5 agents and 250 tasks; the processing times $t_{a,j}$ were drawn uniformly at random. Both variants–with and without fairness constraints–were implemented in Kotlin and solved using CPLEX 22.1. The outer‑approximation algorithm of Section IV-B was realized via CPLEX’s lazy‑constraint callback. Source code for instance generation and both solution methods is available at https://github.com/abhay-singh11/FairTaskAllocation.
@@ V-A Dispersion Control
V-B Price of Fairness
The term “Price of Fairness” (PoF) is defined as the percentage increase in the objective value of the fairness‑constrained MAAP of Section IV-A relative to that of the unconstrained problem (17). This definition is standard in the literature [31, 10]. Formally,
@@ V-B Price of Fairness
Figure 3 illustrates this trend for one representative instance; the other 49 instances show the same qualitative behavior. The increasing ordering $\mathrm{PoF}(\varepsilon,p_{1})\geqslant\mathrm{PoF}(\varepsilon,p_{2})$ for $p_{1}>p_{2}$ reflects the inclusion property of Theorem 2.
V-C Computation Times
@@
VI Conclusion
In this article, we present a countably infinite family of convex models–parameterized by $\varepsilon\in[0,1]$–to control the dispersion of a subset of decision variables in an optimization problem. The parameter $\varepsilon$ bounds dispersion: when $\varepsilon=0$, dispersion is unconstrained; when $\varepsilon=1$, dispersion is forced to zero. This family is indexed by integers $p\geqslant 2$, and all the models in the family are derived from first principles using the fundamental result of equivalence of finite-dimensional norms. For $p=2$ the model is a second‑order cone; for $2<p<\infty$ it is an $\ell_{p}$‑norm cone; and for $p=\infty$ it reduces to a set of linear inequalities.
@@ VI Conclusion
We prove several theoretical properties that align with intuitive notions of fair solutions. We then evaluate the trade‑offs of these models on the MAAP. In particular, the $p=\infty$ variant is appealing both computationally–since the model is linear–and in terms of dispersion control, as increasing $\varepsilon$ yields a more gradual reduction in dispersion. Finally, these models can be integrated into optimization models in domains such as robotics, logistics, and transportation to control dispersion or enforce fairness of a vector of decision variables with minimal modifications to existing solution algorithms.
Acknowledgement

### v3

@@ 4 Multi-Agent Assignment Problem
Although the formulation in (17) is presented as an MILP, its constraint matrix is the node–arc incidence matrix of a bipartite graph, which is totally unimodular. Consequently, the linear programming relaxation has integral extreme points, and solving it via any linear programming solver yields the optimal integer solution schrijver2003combinatorial . Equivalently, one can apply the Hungarian algorithm directly to obtain the optimum in $O(|\mathcal{J}|^{3})$ time kuhn1955hungarian . In contrast, a natural generalization in which each task carries a weight and each agent has a capacity constraint becomes NP‑hard (the generalized assignment problem) cattrysse1992survey . In this work, we restrict attention to the uncapacitated variant in (17), focusing on illustrating how $(\varepsilon,p)$‑fairness models control the dispersion of the processing times $t_{a}$ across agents in the simplest assignment problem, rather than a general variant of the MAAP.
4.1 Fairness-Constrained Multi-Agent Assignment Problem
@@
Although the problem in (17) is solvable in polynomial time, adding (18) makes the resulting problem NP‑complete. Indeed, for two agents with $\varepsilon=1$, it reduces to the classical partition problem korf1998complete that is known to be strongly NP-complete. As noted in Section 2, when $p<\infty$, (18) is a convex constraint—specifically, an $\ell_{p}$‑norm conic constraint (for $p=2$, an $\ell_{2}$‑norm cone is referred to as a second‑order cone). Hence, the fairness‑constrained MAAP is a mixed‑integer convex program for $p<\infty$. When $p=\infty$, applying the reduction in (7) to (18) yields a mixed‑integer linear program that can be solved directly using off‑the‑shelf commercial or open‑source mixed‑integer linear programming solvers. For mixed‑integer convex optimization problems ($2\leqslant p<\infty$), constraint (18) can be lazily outer approximated by linear inequalities duran1986outer , enabling MILP solvers to obtain the solution. In the next section, we briefly present the outer approximation procedure for a general $\ell_{p}$-norm cone. We remark that the outer approximation algorithm is well-understood in the literature, and the presentation in the next section is only intended to make the article self-contained for the readers.
4.2 Outer Approximation of $\ell_{p}$-Norm Cone
We start with the definition of an $\ell_{p}$-norm cone boyd2004convex of dimension $(n+1)$ as
@@ 4.2 Outer Approximation of $\ell_{p}$-Norm Cone
Note that the fairness constraint in (18) for any $p\geqslant 2$ can be represented in the form (19) by introducing an auxiliary variable. It is known in the literature that when convex cones are outer approximated by linear inequalities, it is both theoretically stronger and computationally more efficient to decompose a high-dimensional cone into several lower-dimensional cones and approximate each of those individually, rather than approximating the high-dimensional cone directly lubin2018polyhedral . Converting a high-dimensional conic constraint into many low-dimensional cones is called “cone disaggregation” lubin2018polyhedral . To that end, we present the steps to disaggregate the $(n+1)$-dimensional $\ell_{p}$-norm cone in (19) into $n$ three-dimensional cones plus a single linear constraint:
@@
In the last equivalence, each constraint $Y_{i}^{\tfrac{1}{p}}\,z^{1-\tfrac{1}{p}}\geqslant y_{i}$ is convex for $2\leqslant p<\infty$ and in fact defines a three-dimensional power cone boyd2004convex :
@@
5 Computational Results
This section provides empirical validation of the theoretical results in Section 3 using the formulations from Section 4. All experiments were run on 50 randomly generated instances of the MAAP, each with 5 agents and 250 tasks; the processing times $t_{a,j}$ were drawn uniformly at random. Both variants–with and without fairness constraints–were implemented in Kotlin and solved using CPLEX 22.1. The outer‑approximation algorithm of Section 4.2 was realized via CPLEX’s lazy‑constraint callback. Source code for instance generation and both solution methods is available at https://github.com/abhay-singh11/FairTaskAllocation.
@@ 5.1 Dispersion Control
5.2 Price of Fairness
The term “Price of Fairness” (PoF) is defined as the percentage increase in the objective value of the fairness‑constrained MAAP of Section 4.1 relative to that of the unconstrained problem (17). This definition is standard in the literature bertsimas2011price ; sundar2025parametric . Formally,
@@ 5.2 Price of Fairness
Figure 3 illustrates this trend for one representative instance; the other 49 instances show the same qualitative behavior. The increasing ordering $\mathrm{PoF}(\varepsilon,p_{1})\geqslant\mathrm{PoF}(\varepsilon,p_{2})$ for $p_{1}>p_{2}$ reflects the inclusion property of Theorem 3.2.
5.3 Computation Times
@@
6 Conclusion
@@
The note presents a countably infinite family of convex models–parameterized by $\varepsilon\in[0,1]$–to control the dispersion of a subset of decision variables in an optimization problem. The parameter $\varepsilon$ bounds dispersion: when $\varepsilon=0$, dispersion is unconstrained; when $\varepsilon=1$, dispersion is forced to zero. This family is indexed by integers $p\geqslant 2$, and all the models in the family are derived from first principles using the fundamental result of equivalence of finite-dimensional norms. For $p=2$, the model is a second‑order cone; for $2<p<\infty$, it is an $\ell_{p}$‑norm cone; and for $p=\infty$, it reduces to a set of linear inequalities. The note also establishes several theoretical properties that align with intuitive notions of fair solutions. The $p=\infty$ variant is appealing computationally since the model is linear. We then evaluate the trade‑offs of these models on the MAAP. In particular, the $p=\infty$ variant is appealing both computationally—since the model is linear—and in terms of dispersion control, as increasing $\varepsilon$ yields a more gradual reduction in dispersion. Finally, these models can be integrated into optimization models in many application domains to control dispersion or enforce the fairness of a vector of decision variables with minimal modifications to existing solution algorithms.
Appendix A Proof of Proposition 4
We start with introducing additional notation and a few lemmas that will make the presentation of the proof of the main result cleaner. For any vector $\bm{x}\in\Delta_{n}\setminus\left\{\bm{e}_{1},\dots,\bm{e}_{n},\tfrac{1}{n}\bm{e}\right\}$, we let
$S_{p}=\sum_{i=1}^{n}x_{i}^{p}~\Rightarrow~~S_{p}^{\frac{1}{p}}=\|\bm{x}\|_{p}$ (24)
and
$w_{i}=\frac{x_{i}^{p}}{S_{p}}~~\Rightarrow~~\sum_{i=1}^{n}w_{i}=1\text{ and }\sum_{i=1}^{n}w_{i}^{\frac{1}{p}}=\frac{1}{\|\bm{x}\|_{p}}.$ (25)
Also, applying $\mathcal{NE}(1,p)$ to $\bm{x}$ yields
$\|\bm{x}\|_{p}<1<(D_{p}+1)~\|\bm{x}\|_{p}~~\Longrightarrow~~\|\bm{x}\|_{p}\in\left((D_{p}+1)^{-1},1\right)$ (26)
Observe that $\bm{x}=\bm{e}_{i}$ for any $i\in\{1,\dots,n\}$ if and only if $\|\bm{x}\|_{p}=1$ and $\bm{x}=\tfrac{1}{n}\bm{e}$ if and only if $\|\bm{x}\|_{p}=(D_{p}+1)^{-1}$. Since these values are excluded from the possible values $\bm{x}$ can take, the inequality in (26) is strict.
Lemma 1
For any $\bm{x}\in\Delta_{n}\setminus\left\{\bm{e}_{1},\dots,\bm{e}_{n},\tfrac{1}{n}\bm{e}\right\}$ and integer $p\geqslant 2$, the following inequality holds
$\frac{p}{p-1}\left[\frac{-\ln\|\bm{x}\|_{p}}{1-\|\bm{x}\|_{p}}\right]-\ln n\left[1+\frac{1}{D_{p}}\right]<0$ (27)
Proof
We start the proof by observing that the function
$f(\|\bm{x}\|_{p})=\frac{-\ln\|\bm{x}\|_{p}}{1-\|\bm{x}\|_{p}}~~\text{ strictly decreases for $\|\bm{x}\|_{p}\in(0,1)$.}$
The above result is easy to see by evaluating the derivative and checking that the sign of the derivative is strictly negative. Now, for any fixed $p\geqslant 2$, we define $t_{0}=(D_{p}+1)^{-1}$. Since $\|\bm{x}\|_{p}\mapsto f(\|\bm{x}\|_{p})$ is a strictly decreasing function, for every value of $\|\bm{x}\|_{p}>t_{0}$ we have $f(\|\bm{x}\|_{p})<f(t_{0})$, i.e.,
$-\frac{\ln\|\bm{x}\|_{p}}{1-\|\bm{x}\|_{p}}$ $<-\frac{\ln(D_{p}+1)^{-1}}{1-(D_{p}+1)^{-1}}$
$=\frac{\ln(D_{r}+1)}{1-(D_{r}+1)^{-1}}$
$=\frac{\ln n^{1-\frac{1}{r}}}{1-(D_{r}+1)^{-1}}$
$=\frac{r-1}{r}~\ln n\left[\frac{D_{r}+1}{D_{r}}\right]$
$=\frac{p-1}{p}~\ln n\left[1+\frac{1}{D_{p}}\right]$
$\Rightarrow~~\frac{p}{p-1}\left[\frac{-\ln\|\bm{x}\|_{p}}{1-\|\bm{x}\|_{p}}\right]$ $<\ln n\left[1+\frac{1}{D_{p}}\right]$
$\Rightarrow~~\frac{p}{p-1}\left[\frac{-\ln\|\bm{x}\|_{p}}{1-\|\bm{x}\|_{p}}\right]$ $-\ln n\left[1+\frac{1}{D_{p}}\right]<0.$
We remark that when $\bm{x}=\tfrac{1}{n}\bm{e}$, $\|\bm{x}\|_{p}$ evaluates to $t_{0}$ and the above inequality is satisfied at equality, and the claim of the lemma does not hold. Hence, $\bm{x}\neq\tfrac{1}{n}\bm{e}$ is assumed in the statement of the lemma.
Lemma 2
Suppose $\bm{w}=[w_{1},\dots,w_{n}]^{\intercal}$, then for any $p\geqslant 2$, the function $H(\bm{w})=-\sum_{i}w_{i}\ln w_{i}$ satisfies
$0\leqslant H(\bm{w})\leqslant-\frac{p}{p-1}\ln\|\bm{x}\|_{p}.$ (28)
Proof
The definitions in (25) tell us that $\bm{w}=[w_{1},\dots,w_{n}]^{\intercal}$ is a probability vector. Hence, $H(\bm{w})$ is the Shannon-entropy of $\bm{w}$, which we know is non-negative gallager1968information . To derive the upper bound, we use the definition of Rényi entropy of order $\alpha$ renyi1961measures given below
$H_{\alpha}(\bm{w})=\frac{1}{1-\alpha}\ln\sum_{i}w_{i}^{\alpha}.$ (29)
The following two facts about $H_{\alpha}(\cdot)$ are well known: $H_{1}(\bm{w})=H(\bm{w})$ and $H_{\alpha}(\cdot)$ is a decreasing function of $\alpha$. We now let $\alpha=\frac{1}{p}$. Then since $\alpha<1$ (because $p\geqslant 2$), we have
$H(\bm{w})\leqslant H_{\frac{1}{p}}(\bm{w})=\frac{p}{p-1}\ln\sum_{i=1}^{n}w_{i}^{\frac{1}{p}}=-\frac{p}{p-1}\ln\|\bm{x}\|_{p}.$ (30)
The final equality follows from (25).
Lemma 3
For any integer $p\geqslant 2$ and $\bm{x}\in\Delta_{n}$, given the definitions in (24), (25) and Lemma 2, the following relationship holds
$\ln S_{p}-p\frac{S_{p}^{\prime}}{S_{p}}=H(\bm{w})$
where $S_{p}^{\prime}$ is the derivative of $S_{p}$ with respect to pp.
Proof
We start by computing the derivative of $S_{p}$ with respect to $p$ as
$S_{p}^{\prime}$ $=\sum_{i=1}^{n}x_{i}^{p}\ln x_{i}=S_{p}\sum_{i=1}^{n}w_{i}\ln x_{i}$
$\iff p\frac{S_{p}^{\prime}}{S_{p}}$ $=\sum_{i=1}^{n}w_{i}\ln x_{i}^{p}=\sum_{i=1}^{n}w_{i}\ln(S_{p}w_{i})$
$\iff p\frac{S_{p}^{\prime}}{S_{p}}$ $=\ln S_{p}\left(\sum_{i=1}^{n}w_{i}\right)+\sum_{i=1}^{n}w_{i}\ln w_{i}$
$=\ln S_{p}-H(\bm{w})$
$\iff H(\bm{w})$ $=\ln S_{p}-r\frac{S_{p}^{\prime}}{S_{p}}.$
We now present the proof of Proposition 4 by performing direct sign analysis of the derivative of $F(p)$. We start by taking logarithm of $F(p)$ as
$\ln F(p)\triangleq g(p)=\ln(1-\|\bm{x}\|_{p})-\ln\|\bm{x}\|_{p}-\ln D_{p}.$ (31)
We now let $F^{\prime}(p)$, $g^{\prime}(p)$, $\|\bm{x}\|_{p}^{\prime}$ and $D^{\prime}_{p}$ denote the derivative with respect to $p$ of $F(p)$, $g(p)$, $\|\bm{x}\|_{p}$ and $D_{p}$, respectively. Then, we can differentiate (31), we get
$\frac{F^{\prime}(p)}{F(p)}=g^{\prime}(p)=-\frac{\|\bm{x}\|_{p}^{\prime}}{1-\|\bm{x}\|_{p}}-\frac{\|\bm{x}\|_{p}^{\prime}}{\|\bm{x}\|_{p}}-\frac{D^{\prime}_{p}}{D_{p}}$
Simplifying the above equation further, we obtain
$g^{\prime}(p)$ $=-\frac{\|\bm{x}\|_{p}^{\prime}}{\|\bm{x}\|_{p}(1-\|\bm{x}\|_{p})}-\frac{D^{\prime}_{p}}{D_{p}}$
$=\frac{-1}{(1-\|\bm{x}\|_{p})}\left[\frac{1}{p}\frac{S_{p}^{\prime}}{S_{p}}-\frac{1}{p^{2}}\ln S_{p}\right]-\frac{\ln n}{p^{2}}\left[1+\frac{1}{D_{p}}\right]$
$=\frac{1}{r^{2}}\left[\frac{\ln S_{r}-r\tfrac{S_{r}^{\prime}}{S_{r}}}{1-\|\bm{x}\|_{r}}\right]-\frac{\ln n}{r^{2}}\left[1+\frac{1}{D_{r}}\right]$
[…]
$\leqslant\frac{1}{p^{2}}\left(\frac{p}{p-1}\left[\frac{-\ln\|\bm{x}\|_{p}}{1-\|\bm{x}\|_{p}}\right]-\ln n\left[1+\frac{1}{D_{p}}\right]\right)$
[…]
The penultimate inequality follows from Lemma 2. We have proved that $g^{\prime}(p)<0$. But $F^{\prime}(p)=F(p)\cdot g^{\prime}(p)$ and since $F(p)>0$ for any $\bm{x}$ that satisfies the assumptions of the proposition, we obtain $F^{\prime}(p)<0$. ∎
Acknowledgements.
@@
Conflict of interest
The authors declare that they have no conflict of interest.